OpenStudy (anonymous):

find a number k>0 such that the area bounded by the curves y=X^2 and y=k-x^2 is 72 square units.

OpenStudy (raden):

18. is the first guess :) do u have choices ?

OpenStudy (anonymous):

i dont have any choices., how did you got it?

OpenStudy (raden):

actually, i use by formula : area = (D * sqrt(D))/6a^2 with D =discriminant of quadratic equation and D=b^2-4ac

OpenStudy (raden):

here, known the area bouded of curves y = x^2 and y = k - x^2 first, we take y1 = y2 therefore x^2 = k - x^2 2x^2 - k = 0 a = 2, b = 0, c = -k

OpenStudy (raden):

so, the value of discriminant is D=0^2 - 4(2)(-k) = 8k

OpenStudy (raden):

area = (D * sqrt(D))/6a^2 72 = 8k sqrt(8k) / 6(2)^2 72 = 8k sqrt(8k) / 24 72 = k sqrt(8k) / 3 squaring to both sides, gives us 72^2 = k^2 (8k)/ 9 (9 * 8)^2 = 8k^3/9 9^2 * 8^2 = 8k^3/9 multiply both sides by 9/8, giving us 9^3 * 8 = k^3 9^3 * 2^3 = k^3 (9 * 2) ^3 = k^3 the finally, we get k=18

OpenStudy (raden):

sorry, if my answer is wrong... but this is my opinion :)

OpenStudy (anonymous):

would it have a shorter answer? like use integral?

OpenStudy (anonymous):

its okay :)

OpenStudy (raden):

yeah, this problem actually can solved by integration or by using formula above but it is an easier than by integral :)

OpenStudy (raden):

yeah, maybe this is a shorter answer